# Theory of Microphase Separation in Elastomers

### v1

@@ Abstract
A recent experiment showed elasticity-controlled microphase separation to be a highly effective technique for generating patterned elastomers with complex morphologies [24]. In the experiment, a temperature quench triggers microphase separation in swollen elastomers. Elasticity was observed to arrest phase separation, resulting in a stable bicontinuous microstructure or droplets with a characteristic size of a few microns. This microphase separation plausibly arises because of a pronounced difference in the length scales at which thermodynamics and elasticity operate. This is unlike previous examples, where patterned phases were primarily seen in systems with anisotropic elasticity or external stresses [25] or involving nontrivial phenomena such as cavitation [18, 26].
In this Letter, we introduce a minimal phase-field model that captures the key features of microphase separation in swollen elastomers in the limit of weak segregation. We show that it is possible to address the length-scale discrepancy between elasticity and thermodynamics by employing a continuous order parameter and coarse-graining it further. The resulting long-range interactions lead to the emergence of stable, finite-sized domains whose length scale is governed by the stiffness of the elastomer. Our scaling results for the domain size and microphase separation temperature agree with experimental observations. We also highlight the diversity of the microphase morphologies by constructing a phase diagram and supplementing it with numerical results. Put together, our findings underscore an intricate coupling between thermodynamics and elasticity, opening up novel ways to produce patternable materials for various purposes.
@@
Figure 1: Displacements $\bm{u}(\bm{x})$ occurring below a characteristic length scale $h$ do not stress the elastomer significantly. Using a coarse-grained strain field $\bar{\bm{\varepsilon}}$, such displacements are “blurred away” and filtered out. In our model, we choose $h$ as a multiple $n\xi$ of the end-to-end distance $\xi$ between adjacent cross-links of the polymeric network within the elastomer.
We consider a charge-neutral elastomer consisting of a cross-linked polymer network isotropically swollen with a solvent. Polymer-solvent interaction occurs over typical intermolecular distances (e.g., the size of the solvent molecules). On the other hand, the elastic response of the elastomer stems entirely from the underlying polymer network, which has a much larger, usually mesoscopic, characteristic length scale (Fig. 1). Deformations of the elastomer occurring below this length scale should not engender a significant elastic response. Elastomers can undergo large deformations during swelling, and they are customarily studied using nonlinear elasticity [27]. However, once the elastomer is completely swollen, further elastic deformations are well described using linear elasticity in terms of a three-dimensional (3D) displacement field $\bm{u}(\bm{x})$ defined over points $\bm{x}$ on the elastomer [28]. The resulting strain field is $\bm{\varepsilon}=\frac{1}{2}[\bm{\nabla}\bm{u}+(\bm{\nabla}\bm{u})^{\mathsf{T}}]$, with $(\bm{\nabla}\bm{u})^{\mathsf{T}}$ being the transpose of $\bm{\nabla}\bm{u}$.
@@
Here $K_{h}(\bm{x}-\bm{x})$ is an isotropic, scalar kernel that depends only on the distance $\left|\bm{x}-\bm{x}^{\prime}\right|$ between two points $\bm{x},\bm{x}^{\prime}$ in space. For concreteness, we use a normalized Gaussian kernel $K_{h}({\bm{x}})=(4\pi h^{2})^{-3/2}\mathrm{e}^{-|\bm{x}|^{2}/(4h^{2})}$, with $h$ being a suitable mesoscopic length scale that controls the amount of coarse-graining. Nonetheless, as we demonstrate in the Supplemental Material (SM) [29], our results are independent of our choice for this kernel.
The stress $\bm{\sigma}$ computed using Eq. (1) models the correct elastic response of the elastomer, while simultaneously allowing us to use the strain $\bm{\varepsilon}$ to capture compositional changes at molecular length scales. This model is a particular instance of the Eringen framework [30, 31, 32] of nonlocal elasticity, and it leads to an elastic energy density $w(\bm{\varepsilon})$ of the form
@@
Here we have defined the order parameter (phase field) $\psi(\bm{x})=\phi(\bm{x})-\phi_{\text{c}}$ assuming that the polymer-solvent system has a critical point $(\phi_{\text{c}},T_{\text{c}})$, and take the mixing free-energy density $f(\psi)$ to be in a Landau form
$f(\psi)=\frac{1}{2}a(T-T_{\text{c}})\psi^{2}+\frac{1}{4}b\psi^{4},$ (5)
with $a$ and $b$ being positive phenomenological constants. Also included in Eq. (4) is the elastic energy density $w(\bm{\varepsilon})$ and a gradient-squared term with an interfacial parameter $\kappa>0$ to penalize spatial variations in $\psi$. Finally, $\eta$ is a Lagrange multiplier to constrain the mean value of $\psi$ to $\psi_{0}=\phi_{0}-\phi_{\text{c}}$, thereby conserving the total volume of the polymer network.
@@
$\operatorname{tr}\bm{\varepsilon}=\bm{\nabla\cdot}\bm{u}\approx-\phi_{\text{c}}^{-1}\psi.$ (6)
Compositional changes in the polymer volume fraction during temperature quenches arise primarily via solvent diffusion. This allows us to disregard shear deformations and use Eq. (6) to write the total elastic energy as a binary interaction in $\psi$ mediated by the coarse-graining kernel $K_{h}$.
For linear stability analysis of Eq. (4), we express the order parameter ${\psi}(\bm{x})$ and the kernel $K_{h}(\bm{x})$ in terms of their Fourier transforms, $\psi_{\bm{q}}=\int\mathrm{d}^{3}x\,\mathrm{e}^{-i\bm{q}\cdot\bm{x}}\,\psi(\bm{x})$ and $K_{h}(\bm{q})=\mathrm{e}^{-h^{2}q^{2}}$. The total free energy in Fourier space then takes the form (SM [29])
@@
$F_{\bm{q}}=a(T-T_{\text{c}})+\kappa q^{2}+M\mathrm{e}^{-h^{2}q^{2}}.$ (8)
Here $q=\left|\bm{q}\right|$ and $M=(\lambda+2\mu)/\phi_{\text{c}}^{2}$ is the rescaled longitudinal modulus [33, 34] of the swollen elastomer. Also, as we are limiting ourselves to linear analysis, we have left the nonquadratic terms of Eq. (7) untransformed.
The second term in Eq. (8), which favors long-range (small $q$) modulations in $\psi$, measures the energy cost to create interfaces. Meanwhile, the elastic term $M\mathrm{e}^{-h^{2}q^{2}}$ favors short-range (large $q$) modulations. Hence, we expect the emergence of a stable, spatially modulated phase at an intermediate length scale, provided that the elastic term is adequately large. The characteristic size of the modulated phase scales as $\Lambda\sim 2\pi q_{\text{m}}^{-1}$, where $q_{\text{m}}$ is the critical wavenumber at which $F_{\bm{q}}$ acquires its minimum. From Eq. (8), we see that $F_{\bm{q}}$ has a minimum at a nonzero $q_{\text{m}}$ given by
@@
only if the dimensionless parameter $\gamma=Mh^{2}/\kappa>1$. The parameter $\gamma$, which measures the relative importance of elastic and interfacial effects, is analogous to the (inverse) elastocapillary number [35, 20]. If the elastic energy cost exceeds the cost to form interfaces ($\gamma>1$), the system can minimize its total energy by creating many stable, finite-sized domains, resulting in microphase separation.
During a temperature quench from the uniform phase with $\psi(\bm{x})=\psi_{0}$, the onset of microphase separation is indicated by linear instability in the fluctuations. By setting $\psi(\bm{x})=\psi_{0}+\delta\psi(\bm{x})$ and expanding Eq. (7) to quadratic order in the fluctuations $\delta\psi(\bm{x})$, we find that such an instability occurs at a temperature where $F_{\bm{q}}=-3b\psi_{0}^{2}$ and $q=q_{\text{m}}$. This leads to an estimate for the temperature $T_{\text{m}}$ at which microphase separation sets in, which we find to be
$T_{\text{m}}(\psi_{0})=T_{\text{c}}-a^{-1}\left[3b\psi_{0}^{2}+M\gamma^{-1}\left(1+\ln\gamma\right)\right].$ (10)
Because $\gamma>1$, we see that $T_{\text{m}}(\psi_{0})<T_{\text{c}}$ for all $\psi_{0}$. Furthermore, $T_{\text{m}}$ decreases linearly with the modulus $M$, showing that deeper temperature quenches are required to induce microphase separation in stiffer elastomers.
@@
Apart from the intermolecular distance, a relevant length scale in swollen elastomers is the root-mean-square end-to-end distance $\xi$ of the segments between adjacent cross-links in the polymer network [36, 37, 38, 39] (Fig. 1). Taking each segment to be a freely-jointed chain with a Flory ratio $C_{\infty}$ [40], composed of $N$ repeat units of length $\ell$, we have $\xi=\phi_{\text{c}}^{-1/3}(C_{\infty}N)^{1/2}\ell$ [40]. Here, the additional factor $\phi_{\text{c}}^{-1/3}$ accounts for the stretch due to isotropic swelling. If the segments and the repeat units have molecular masses $m_{\text{s}}$ and $m_{\text{r}}$, respectively, then $N=m_{\text{s}}/m_{\text{r}}$. Assuming that the elastomer has a mass density $\rho$, its Young’s modulus in the dry state is $Y=3\rho k_{\text{B}}T/m_{\text{s}}$ [41]. Using this expression to write $m_{\text{s}}$ and $N$ in terms of $Y$, we find that the end-to-end distance scales as [37, 38]
$\xi\sim\phi_{\text{c}}^{-1/3}(3B/Y)^{1/2},$ (11)
where the material-dependent parameter $B$ is given by
$B=C_{\infty}\rho\ell^{2}k_{\text{B}}T/m_{\text{r}},$ (12)
@@
In order to estimate the domain size $\Lambda$ of the microphases, we note that the rescaled longitudinal modulus $M$ of an elastomer is related to its dry Young’s modulus $Y$ via $M\sim\frac{1}{3}\phi_{\text{c}}^{-5/3}Y$ [42, 29]. Interface formation occurs at intermolecular length scales, so we estimate the interfacial parameter as $\kappa\sim k_{\text{B}}T/\ell$ [43]. With the choice $h=n\xi$, the parameter $\gamma$ is independent of $Y$, and using Eqs. (9) and (11) we find the scaling
$\Lambda\sim 2\pi\left[\frac{3Bn^{2}\phi_{\text{c}}^{-2/3}}{Y\ln\left(Bn^{2}\kappa^{-1}\phi_{\text{c}}^{-7/3}\right)}\right]^{1/2}.$ (13)
In Fig. 2(a), we compare the experimental results and Eq. (13) and find good agreement between the two. Furthermore, as we see from Fig. 2(b), the microphase separation temperature $T_{\text{m}}$ linearly decreases with $Y$, which is consistent with the prediction in Eq. (10). We show in the SM [29] that these scalings for $\Lambda$ and $T_{\text{m}}$ are agnostic to the choice of the kernel $K_{h}$ in Eq. (2).
@@
Close to the critical point, the domain boundaries of the microphases are diffuse (weak segregation), and they are well approximated as modulations in the order parameter $\psi$ with a wavenumber $q=q_{\text{m}}$ [44, 45]. A phase diagram in the $(\phi_{0},T)$ plane constructed using this one-mode approximation is presented in Fig. 3(a), with the analytical steps detailed in the SM [29]. For simplicity, we have only examined 2D modulations in the phase diagram. Nonetheless, it shows excellent agreement with the equilibrium phases found by numerically minimizing the free energy in 3D [Figs. 3(b) and 3(c)]. Near the critical point, there are three distinct phases: a uniform phase, a droplet (hexagonal) phase consisting of solvent-rich droplets embedded within the elastomer, and a stripe phase composed of alternating solvent-rich and solvent-deficient layers. An “inverted” droplet phase also appears at low $\phi_{0}$. The phase diagram has the same topology as phase diagrams for other modulated systems [46, 47, 48, 49] and, in principle, there should be regions of phase coexistence. However, for the experimental parameter ranges used here, the widths of these regions are very small, and therefore are not depicted.
@@
The phase diagram in Fig. 3 also shows good agreement with the experimental results and predicts the onset of microphase separation well. Experimentally, droplets are seen in soft elastomers with $Y\lesssim 40{\kern 2.333pt}\text{kPa}$. Only bicontinuous structures (different from stripes and droplets) are observed in stiff elastomers. However, because of the generic topology of the theoretical phase diagram, irrespective of the stiffness, we expect the droplet phase to always appear first during an off-critical temperature quench. This suggests that some other mechanism is responsible for the emergence of the bicontinuous structures in stiffer samples, e.g., shear deformations or nonlinear effects, which we have neglected.
@@
Using a phase-field model for swollen elastomers, we have predicted the possibility of a microphase separation arising from an imbalance between the intermolecular length scale and the mesoscopic coarseness of network elasticity. The elastomer remains stable with an intrinsically selected length scale if the free-energy contribution from the network elasticity is adequately large compared to the interfacial energy costs. Our scaling predictions for the domain size and the microphase separation temperature as a function of the elastic moduli are consistent with recent experimental observations [24].
In a recent theoretical work [50] inspired by the same experiments, a 1D nonlocal model was used to derive the scaling $\Lambda\sim Y^{-1/2}h^{1/2}\kappa^{1/4}$ in the strong-segregation limit, taking the nonlocality scale $h$ and the Young’s modulus $Y$ to be independent. This scaling is different from our 3D result in Eq. (13) for weak segregation, which also takes into account the inter-dependence of $h$ and $Y$. Extensions of our theory to ternary systems in the strong segregation regime could help elucidate the non-power-law scaling of the domain size observed in earlier experiments [12]. Our framework can also be used to study other systems where elasticity and thermodynamics often operate at different length scales, such as certain porous materials [51, 52], metals [53, 2], colloidal suspensions [54], etc.
@@
Figure 3: (a) Phase diagram in the $(\phi_{0},T)$ plane for an elastomer with a dry Young’s modulus $Y=800{\kern 2.333pt}\text{kPa}$. Here $T$ is the temperature, and $\phi_{0}$ is the mean polymer volume fraction. Other parameters are the same as in Fig. 2. The solid curves show the phase boundaries (binodals). Phase coexistence regions are not depicted as they are very narrow. The dashed curve depicts the microphase separation temperature $T_{\text{m}}$ from Eq. (10) with a shifted critical temperature $T_{\text{c}}^{\prime}=T_{\text{m}}(0)$. The open circles represent experimental results from Ref. [24]. (b), (c) Equilibrium morphologies of the elastomer obtained by numerically minimizing the free energy, Eq. (7), with the corresponding $(\phi_{0},T)$ values marked in (a). Solvent-rich (small $\phi_{0}$) and solvent-deficient (large $\phi_{0}$) regions are highlighted in red and blue, respectively.
@@
We thank Ram Adar, Amit Kumar, and Sam Safran for useful conversations. M.M. acknowledges partial support through the Bloomfield Fellows Program at Tel Aviv University. H.D. thanks Lev Truskinovsky for helpful discussions and acknowledges support from the Israel Science Foundation (ISF Grant No. 1611/24). D.A. acknowledges support from the Israel Science Foundation (ISF Grant No. 226/24).
@@
Manu Mannattil,1,2 Haim Diamant,2 David Andelman2
@@
This Supplemental Material is organized as follows: In Sec. I, we derive an expression for the free energy of swollen elastomers. A comparison of recent experimental results with our model is presented in Sec. II, with phase diagrams discussed in Sec. III. In Sec. IV, we describe a general coarse-graining procedure and prove that the domain size scaling is independent of the specific choice of the coarse-graining kernel. Finally, in Sec. V, we discuss the numerical techniques used in this work.

### v2

@@ Abstract
A recent experiment showed elasticity-controlled microphase separation to be a highly effective technique for generating patterned elastomers with complex morphologies [24]. In the study, a temperature quench is used to trigger microphase separation in elastomers swollen with a solvent. The results are reminiscent of older observations of phase separation and critical density fluctuations in swollen gels as the temperature is lowered [25, 26, 27, 28]. In the new experiments, however, the elastomer does not fully phase separate from the solvent, and instead forms stable bicontinuous microstructures or droplets whose sizes are determined by the stiffness of the elastomer. This microphase separation plausibly arises because of a pronounced difference in the length scales at which thermodynamics and elasticity operate [24]. This is unlike previous examples, where patterned phases were primarily seen in systems with anisotropic elasticity or external stresses [29] or involving nontrivial phenomena such as cavitation [18, 30].
In this Letter, we introduce a phase-field model that captures the key features of microphase separation in swollen elastomers in the limit of weak segregation. Recent theoretical work [31], also inspired by the aforementioned experiments, has demonstrated that the length-scale discrepancy between elasticity and thermodynamics in elastomers can be resolved using nonlocal theories of elasticity [32, 33, 34]. Nonlocal approaches have also been employed in other systems with scale-dependent phenomena, such as certain porous materials [35] and DNA elasticity [36].
The stiffness of elastomers arise primarily due to strain-induced changes in the configurational entropy of polymer chains [37, 38, 39]. Our scaling results for the domain size and microphase separation temperature, obtained by using results from rubber (entropic) elasticity and incorporating nonlocal effects, agree with the experimental observations. We also highlight the diversity of the microphase morphologies by constructing a phase diagram and supplementing it with numerical results. Put together, our findings underscore an intricate coupling between thermodynamics and elasticity, opening up novel ways to produce patternable materials for various purposes.
@@
Figure 1: Displacements $\bm{u}(\bm{x})$ occurring below a characteristic length scale $h$ do not stress the elastomer significantly. Using a coarse-grained strain field $\bar{\bm{\varepsilon}}$, such displacements are “blurred away” and filtered out. In our model, we choose $h$ as a multiple $n\xi$ of the end-to-end distance $\xi$ between adjacent crosslinks of the polymeric network within the elastomer.
We consider a charge-neutral elastomer consisting of a crosslinked polymer network isotropically swollen with a solvent. Polymer-solvent interaction occurs over typical intermolecular distances (e.g., the size of the solvent molecules). On the other hand, the elastic response of the elastomer stems entirely from the underlying polymer network, which has a much larger, usually mesoscopic, characteristic length scale (Fig. 1). Deformations of the elastomer occurring below this length scale should not engender a significant elastic response. Elastomers can undergo large deformations during swelling, and they are customarily studied using nonlinear elasticity [40]. However, once the elastomer is completely swollen, further elastic deformations are well described using linear elasticity in terms of a three-dimensional (3D) displacement field $\bm{u}(\bm{x})$ defined over points $\bm{x}$ on the elastomer [41]. The resulting strain field is $\bm{\varepsilon}=\frac{1}{2}[\bm{\nabla}\bm{u}+(\bm{\nabla}\bm{u})^{\mathsf{T}}]$, with $(\bm{\nabla}\bm{u})^{\mathsf{T}}$ being the transpose of $\bm{\nabla}\bm{u}$.
@@
Here $K_{h}(\bm{x}-\bm{x}^{\prime})$ is an isotropic, scalar kernel that depends only on the distance $\left|\bm{x}-\bm{x}^{\prime}\right|$ between two points $\bm{x},\bm{x}^{\prime}$ in space. For concreteness, we use a normalized Gaussian kernel $K_{h}({\bm{x}})=(4\pi h^{2})^{-3/2}\mathrm{e}^{-|\bm{x}|^{2}/(4h^{2})}$, with $h$ being a suitable mesoscopic length scale that controls the extent of coarse-graining. Nonetheless, as we demonstrate in the Supplemental Material (SM) [42], our results are independent of our choice for this kernel.
The stress $\bm{\sigma}$ computed using Eqs. (1) and (2) models the correct elastic response of the elastomer, while simultaneously allowing us to use the strain $\bm{\varepsilon}$ to capture compositional changes at molecular length scales. This model is a particular instance of the Eringen framework [32, 33, 34] of nonlocal elasticity, and it leads to an elastic energy density $w(\bm{\varepsilon})$ of the form
@@
Here we have defined the order parameter (phase field) $\psi(\bm{x})=\phi(\bm{x})-\phi_{*}$ assuming that the homogeneous system has a “critical” point $(\phi_{*},T_{*})$, and take the free-energy density $f(\psi)$ to be in a Landau form
$f(\psi)=\frac{1}{2}a(T-T_{*})\psi^{2}+\frac{1}{4}b\psi^{4},$ (5)
with $a$ and $b$ being positive phenomenological constants. Contributions from polymer-solvent mixing are included in this phenomenological $f(\psi)$, with the $\psi^{4}$ term playing an additional role in stabilizing phase separation. For polymer networks cross-linked in solution, the critical temperature $T_{*}$ is likely to be close to the theta temperature before cross-linking [69]. Similar free-energy densities have been used to model swelling and deswelling of gels [70, 71, 72, 73]. Also included in Eq. (4) is the elastic energy density $w(\bm{\varepsilon})$ and a gradient-squared term with an interfacial parameter $\kappa>0$ to penalize spatial variations in $\psi$. Finally, $\eta$ is a Lagrange multiplier to constrain the mean value of $\psi$ to $\psi_{0}=\phi_{0}-\phi_{*}$, thereby conserving the total volume of the polymer network.
@@
$\operatorname{tr}\bm{\varepsilon}=\bm{\nabla\cdot}\bm{u}\approx-\phi_{*}^{-1}\psi.$ (6)
Compositional changes in the polymer volume fraction during temperature quenches arise primarily via solvent diffusion. This allows us to disregard shear deformations and use Eqs. (3) and (6) to write the total elastic energy as a binary interaction in $\psi$ mediated by the coarse-graining kernel $K_{h}$.
For linear stability analysis of Eq. (4), we express the order parameter ${\psi}(\bm{x})$ and the kernel $K_{h}(\bm{x})$ in terms of their Fourier transforms, $\psi_{\bm{q}}=\int\mathrm{d}^{3}x\,\mathrm{e}^{-i\bm{q}\cdot\bm{x}}\,\psi(\bm{x})$ and $K_{h}(\bm{q})=\mathrm{e}^{-h^{2}q^{2}}$. Upon expressing the quadratic part of the total free energy in Fourier space, we find (SM [42])
@@
$F_{\bm{q}}=a(T-T_{*})+\kappa q^{2}+M\mathrm{e}^{-h^{2}q^{2}}.$ (8)
Here $q=\left|\bm{q}\right|$ and $M=(\lambda+2\mu)/\phi_{*}^{2}$ is the rescaled longitudinal modulus [74, 75] of the swollen elastomer.
The second term in Eq. (8), which favors long-range (small $q$) modulations in $\psi$, measures the energy cost to create interfaces. Meanwhile, the elastic term $M\mathrm{e}^{-h^{2}q^{2}}$ favors short-range (large $q$) modulations. Hence, we expect the emergence of a stable, spatially modulated phase at an intermediate length scale, provided that the elastic term is adequately large. The characteristic size of the modulated phase scales as $\Lambda\sim 2\pi q_{\text{m}}^{-1}$, where $q_{\text{m}}$ is the wavenumber at which $F_{\bm{q}}$ acquires its minimum.
From Eq. (8), we see that $F_{\bm{q}}$ has a minimum at a nonzero $q_{\text{m}}$ given by
@@
only if the dimensionless parameter $\gamma=Mh^{2}/\kappa>1$. The parameter $\gamma$, which measures the relative importance of elastic and interfacial effects, is analogous to the (inverse) elastocapillary number [76, 20] and the Lifshitz point [77] of Eq. (7) is located at $\gamma=1$ (SM [42]). If the elastic energy cost exceeds the cost to form interfaces ($\gamma>1$), the system can minimize its total energy by creating many stable, finite-sized domains, resulting in microphase separation. Note that if $h=0$ in Eq. (2) and the system exhibits a local elastic response, we can recover known results for swollen polymer networks from the free energy in Eq. (7), such as the onset of spinodal decomposition at temperatures where the osmotic longitudinal modulus vanishes, negative Poisson’s ratio etc. [78] (SM [42]).
During a temperature quench from the uniform phase with $\psi(\bm{x})=\psi_{0}$, the onset of microphase separation is indicated by linear instability in the order-parameter fluctuations. Upon expressing $\psi(\bm{x})=\psi_{0}+\delta\psi(\bm{x})$ and expanding the free energy in Eq. (7) up to $\mathcal{O}(\delta\psi^{2})$ in the fluctuations $\delta\psi(\bm{x})$, we determine that the instability arises at a temperature where $F_{\bm{q}}=-3b\psi_{0}^{2}$ and $q=q_{\text{m}}$. This provides an estimate for the temperature $T_{\text{micro}}$ at which microphase separation begins, which we find to be
$T_{\text{micro}}(\psi_{0})=T_{*}-a^{-1}\left[3b\psi_{0}^{2}+M\gamma^{-1}\left(1+\ln\gamma\right)\right].$ (10)
Clearly, $T_{\text{micro}}$ decreases linearly with the modulus $M$, showing that deeper temperature quenches are required to induce microphase separation in stiffer elastomers.
@@
Polydimethylsiloxane (PDMS) elastomers, such as the ones used in the experiments in Ref. [24], are susceptible to chain entanglement effects that can alter their elastic response substantially, particularly at low crosslink densities [79, 80, 81, 82]. However, based on the observed variation of the Young’s modulus with crosslink density (detailed in the SM [42]), we judge entanglement effects to be negligible, enabling us to use classical rubber elasticity theory in our analysis.
Apart from the intermolecular distance, a relevant length scale in elastomers is the root-mean-square end-to-end distance $\xi$ of the strands between adjacent crosslinks in the polymer network [83, 84, 85, 86] (Fig. 1). Taking each strand to be a freely-jointed chain with a Flory ratio $C_{\infty}$ [37], composed of $N$ repeat units of length $\ell$, we have $\xi^{2}\sim\frac{1}{2}C_{\infty}N\ell^{2}$ [38, 86]. Here the factor of $\frac{1}{2}$ is an estimate assuming the network junctions have tetrafunctional connectivity. If the strands and the repeat units have molecular masses $m_{\text{s}}$ and $m_{\text{r}}$, respectively, then $N=m_{\text{s}}/m_{\text{r}}$. Assuming that the elastomer has a mass density $\rho$, its Young’s modulus in the dry state is $Y=3\rho k_{\text{B}}T/2m_{\text{s}}$ [39]. Using this expression to write $m_{\text{s}}$ and $N$ in terms of $Y$, we find that the end-to-end distance scales as [84, 85]
$\xi\sim(3B/Y)^{1/2},$ (11)
@@
In order to estimate the domain size $\Lambda$ of the microphases, we note that the rescaled longitudinal modulus $M$ of a swollen elastomer is related to its dry Young’s modulus $Y$ via $M\sim\frac{1}{3}\phi_{*}^{-5/3}Y$ [28, 42]. Interface formation occurs at intermolecular length scales, so we estimate the interfacial parameter as $\kappa\sim k_{\text{B}}T/\ell$ [87]. With the choice $h=n\xi$, the parameter $\gamma$ is independent of $Y$, and using Eqs. (9) and (11) we find the scaling
$\Lambda\sim 2\pi\left[\frac{3Bn^{2}}{Y\ln\left(Bn^{2}\kappa^{-1}\phi_{*}^{-5/3}\right)}\right]^{1/2}\hskip-10.00002pt.$ (12)
The scaling $\Lambda\sim Y^{-1/2}$ above is markedly different from what one would expect on dimensional grounds alone (the so-called rheological mesh size of polymer networks that scales as $(k_{\text{B}}T/Y)^{1/3}$ [88, 89]). In Fig. 2(a), we compare the experimental results and Eq. (12) and find good agreement between the two. Furthermore, as we see from Fig. 2(b), the microphase separation temperature $T_{\text{micro}}$ linearly decreases with $Y$, which is consistent with the prediction of Eq. (10). Using the $T_{\text{micro}}$ data, one can estimate the parameters $(\phi_{*},T_{*})$ appearing in Eq. (5). We show in the SM [42] that these scalings for $\Lambda$ and $T_{\text{micro}}$ are agnostic to the choice of the kernel $K_{h}$ in Eq. (2).
@@
Close to the critical point, the microphase domain boundaries are diffuse (weak segregation), and they are well approximated as modulations in the order parameter $\psi$ with a wavenumber $q=q_{\text{m}}$ [90, 91]. A phase diagram in the $(\phi_{0},T)$ plane constructed using this one-mode approximation is presented in Fig. 3(a), with the analytical steps detailed in the SM [42]. For simplicity, we have only examined 2D modulations in the phase diagram. Nonetheless, it shows excellent agreement with the equilibrium phases found by numerically minimizing the free energy in 3D [Figs. 3(b) and 3(c)]. Near the critical point, there are three distinct phases: a uniform phase, a droplet (hexagonal) phase consisting of solvent-rich droplets embedded within the elastomer, and a stripe phase composed of alternating solvent-rich and solvent-deficient layers. An “inverted” droplet phase also appears at low $\phi_{0}$.
In Fig. 3(a), the first-order phase-transition curves that divide the different phases converge at a critical point $T_{*}^{\prime}$, where a second-order transition between the uniform and the stripe phase is possible. The phase diagram has the same topology as phase diagrams for block copolymers [92, 93] and other systems displaying modulated phases [94, 95, 96, 97], which are often characterized by a Landau–Brazovskii free energy. We show in the SM [42] that the free energy in Eq. (7) can be simplified to this form, explaining the generic nature of the phase diagram, which also has regions of phase coexistence [42]. However, for the experimental parameter ranges used here, the widths of these regions are very small, and therefore are not depicted. The absence of substantial regions of phase coexistence may account for the apparent lack of hysteresis seen in the experiments [24].
@@
The phase diagram in Fig. 3(a) shows good agreement with the experimental results and predicts the onset of microphase separation well. Experimentally, droplets are seen in soft elastomers with $Y\lesssim 40{\kern 2.333pt}\text{kPa}$. Only bicontinuous structures (different from stripes and droplets) are observed in stiffer elastomers. However, because of the generic topology of the theoretical phase diagram, irrespective of the stiffness, we expect the droplet phase to always appear first during an off-critical temperature quench. This suggests that some other mechanism is responsible for the emergence of bicontinuous structures in stiffer samples, e.g., shear deformations or nonlinear effects, which we have neglected. Further consistency with experiments is seen upon examining the static structure factor, found using Eq. (7) as $S(q)\sim F_{\bm{q}}^{-1}$. It peaks at $q=q_{\text{m}}$, given in Eq. (9), and explains the smooth increase of the scattering intensity at a fixed $q$ during a temperature quench as seen in the experiments (SM [42]).
@@
Figure 3: (a) Phase diagram in the $(\phi_{0},T)$ plane for an elastomer with a dry Young’s modulus $Y=800{\kern 2.333pt}\text{kPa}$. Here $T$ is the temperature, and $\phi_{0}$ is the mean polymer volume fraction. Other parameters are the same as in Fig. 2. The solid curves show the phase boundaries (binodals). Phase coexistence regions are not depicted as they are very narrow. The dashed curve depicts the microphase separation temperature $T_{\text{micro}}(\psi)$ from Eq. (10) with a shifted critical temperature $T_{*}^{\prime}=T_{\text{micro}}(0)$. The open circles represent experimental results from Ref. [24]. (b), (c) Equilibrium morphologies of the elastomer obtained by numerically minimizing the free energy, Eq. (7), with the corresponding $(\phi_{0},T)$ values marked in (a). Solvent-rich ($\phi<\phi_{0}$) and solvent-deficient ($\phi>\phi_{0}$) regions are highlighted in red and blue, respectively.
Using a phase-field model for swollen elastomers, we have predicted the possibility of a microphase separation arising from an imbalance between the intermolecular length scale and the mesoscopic coarseness of network elasticity. The elastomer remains stable with an intrinsically selected length scale if the free-energy contribution from the network elasticity is adequately large compared to the interfacial energy costs. Our scaling predictions for the domain size $\Lambda$ and the microphase separation temperature $T_{\text{micro}}$ as a function of the elastic moduli are consistent with recent experimental observations [24].
As the number of repeat units $N$ between the crosslinks follows the scaling $N\sim Y^{-1}$, we find $\Lambda\sim N^{1/2}$ and a linear dependence between $T_{\text{micro}}$ and $N^{-1}$. Intriguingly, similar scaling behaviors have been experimentally observed in crosslinked polymer blends [98, 99] and were predicted earlier by de Gennes [100] using a phenomenological model that draws an analogy to electrostatics. The similarity in the scaling suggests that the internal elastic response of these blends may be nonlocal.
Our use of nonlocal elasticity was motivated by a recent theoretical study [31] inspired by the same experiments on elastomers. In this study, a one-dimensional (1D) nonlocal model was used to obtain the scaling $\Lambda\sim Y^{-1/2}h^{1/2}\kappa^{1/4}$ in the strong-segregation limit, taking the nonlocality scale $h$ and the Young’s modulus $Y$ to be independent. This scaling is different from our 3D result for weak segregation, Eq. (12), which also takes into account the inter-dependence of $h$ and $Y$, incorporating results from rubber elasticity. Furthermore, our model predicts a first-order transition from the uniform phase to various patterned phases. Conversely, in the 1D nonlocal model, a line of second-order transitions was predicted for large $Y$ based on detailed numerical analyses [31]. Differences in the dimensionalities of the two models may explain this discrepancy (SM [42]).
Microphase separation in elastomers closely resembles that in block copolymers and other systems exhibiting modulated phases [90]. As we discuss in the SM [42], we expect it to be a rich source of related phenomena such as fluctuation-induced first-order transitions [101, 93], Lifshitz behavior [102, 103], microemulsion phases, etc. [104, 105]. Other experimentally relevant theoretical questions include the kinetics of phase separation [70, 106], effect of quenched impurities and network heterogeneities [107, 108, 109], volume phase transitions [110, 111], etc. Finally, extensions of our theory to ternary systems in the strong segregation regime could help elucidate the non-power-law scaling of the domain size observed in earlier experiments [12].
@@
We thank Ram Adar, Amit Kumar, Yitzhak Rabin, Sam Safran, and Lev Truskinovsky for useful conversations. M.M. acknowledges partial support through the Bloomfield Fellows Program at Tel Aviv University. H.D. acknowledges support from the Israel Science Foundation (ISF Grant No. 1611/24). D.A. acknowledges support from the Israel Science Foundation (ISF Grant No. 226/24).
@@
Manu Mannattil,1,2,3 Haim Diamant,1,3 and David Andelman2,3
@@
3Center for Physics and Chemistry of Living Systems, Tel Aviv University, Tel Aviv 69978, Israel
This Supplemental Material is organized as follows: In Sec. I, we derive an expression for the free energy of swollen elastomers. A comparison of recent experimental results with our model is presented in Sec. II. Phase diagrams are discussed in Sec. III, where we also compare our model to other pattern-forming systems. The effects of fluctuations, scattering properties, and Lifshitz behavior are presented in Sec. IV. In Sec. V, we describe a general coarse-graining procedure and prove that the domain size scaling is independent of the specific choice of the coarse-graining kernel. Finally, in Sec. VI, we discuss the numerical techniques used in this work.